The perimeter of a rectangular swimming pool is 154m.its length is 2 m more than twice its breadth.whatare the length and the breadth of the pool?
Answer by rfer(16322)

2x+2(2x+2)=154
2x+4x+4=154
6x=154-4
x=150/6
x=25m
2x+2=52m